What volume of carbon dioxide ( at S.T.P.) is produced during the complete combustion of 200g of methane with a 90% Purity?
1
Expert's answer
2020-05-19T08:59:20-0400

At first we have to calculate mass and amount of pure methane (CH4):

m(pure CH4) = 200 g *0.9 = 180 g - mass of pure methane.

n(pure CH4) = 180 g/16.04 g/mol = 11.22 mol - amount of pure methane

Then, according to reaction:

CH4 + 2O2 = CO2 + 2H2O

We can find out that amount of carbon dioxide produced will be equal to amount of pure methane (because coefficients in equation before them is 1). Therefore n(CO2) = 11.22 mol.

Finally,we need to calculate volume of CO2:

V(CO2) = n(CO2)*Vmolar, Vmolar of gases is known and equal to 22.4 L/mol

So, V(CO2) = 11.22 mol * 22.4 L/mol = 251.33 L

Answer: Volume of carbon dioxide is 251.33 L.